OpenStudy (anonymous):

HELPPPPPPPPPPPP f(x)= x-2 g(x)= x^2-11x+18 (f+g)(x) add the functions

OpenStudy (aum):

(f+g)(x) = f(x) + g(x) = x - 2 + x^2 - 11x + 18 = ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

i dont know how to solve it

OpenStudy (aum):

You just need to combine like terms: (f+g)(x) = f(x) + g(x) = x - 2 + x^2 - 11x + 18 = x^2 - 10x + 16 That is all. combining like terms means: x - 11x = -10x -2 + 18 = 16 Answer: (f+g)(x) = x^2 - 10x + 16